### v3

# Mitigating Biases of Large Language Models in Stance Detection with Counterfactual Augmented Calibration

Abstract
Stance detection is critical for understanding the underlying position or attitude expressed toward a topic. Large language models (LLMs) have demonstrated significant advancements across various natural language processing tasks including stance detection, however, their performance in stance detection is limited by biases and spurious correlations inherent due to their data-driven nature. Our statistical experiment reveals that LLMs are prone to generate biased stances due to sentiment-stance spurious correlations and preference towards certain individuals and topics. Furthermore, the results demonstrate a strong negative correlation between stance bias and stance detection performance, underscoring the importance of mitigating bias to enhance the utility of LLMs in stance detection. Therefore, in this paper, we propose a Counterfactual Augmented Calibration Network (FACTUAL), which a novel calibration network is devised to calibrate potential bias in the stance prediction of LLMs. Further, to address the challenge of effectively learning bias representations and the difficulty in the generalizability of debiasing, we construct counterfactual augmented data. This approach enhances the calibration network, facilitating the debiasing and out-of-domain generalization. Experimental results on in-target and zero-shot stance detection tasks show that the proposed FACTUAL can effectively mitigate biases of LLMs, achieving state-of-the-art results.
Mitigating Biases of Large Language Models in Stance Detection with Counterfactual Augmented Calibration
Ang Li1,3,4∗, Jingqian Zhao1,3thanks:  The first two authors contribute equally to this work., Bin Liang1,3,4, Lin Gui5, Hui Wang2, Xi Zeng6, Xingwei Liang1,3, Kam-Fai Wong4, and Ruifeng Xu1,2,3thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China 2 Peng Cheng Laboratory, Shenzhen, China 3 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 4 The Chinese University of Hong Kong, Hong Kong, China  5 King’s College London, UK 6 The 30th Research Institute of China Electronics Technology Group Corporation, Chengdu, China {angli,zhaojingqian}@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
1 Introduction
Stance detection aims at automatically identifying the author’s opinionated standpoint or attitude (e.g., Favor, Against, or Neutral) expressed in the content towards a specific target, topic, or proposition Somasundaran and Wiebe (2010); Mohammad et al. (2016). With the development of social media platforms, stance detection plays a pivotal role in analyzing public opinion on social media topics Jang and Allan (2018); Ghosh et al. (2019); Stefanov et al. (2020); Sun et al. (2018); Chen et al. (2021).
Large Language Models (LLMs), such as ChatGPT [^1] https://openai.com/blog/chatgpt/, Bard [^2] https://bard.google.com/, and LLaMA Touvron et al. (2023), have demonstrated impressive language comprehension and task-handling capabilities by leveraging extensive corpus and knowledge. However, their data-driven nature makes them susceptible to biases and spurious correlations embedded in pre-training data. In stance detection, which requires interpreting the relationship between a sentence and a specific topic, clues from any isolated aspects could become spurious and lead to biased stances.
Figure 1: An example demonstrates two types of biases encountered by large language models in stance detection tasks (shown at the top and bottom) as well as unbiased stance rationale (depicted in the middle).
Our experiment reveals that LLMs are influenced by two types of bias patterns in stance detection: 1) sentiment-stance spurious correlations, where the sentiment expressed in the sentence misleads the stance judgment towards a particular target, and 2) target preference bias, where LLMs exhibit a certain stance bias towards certain individuals or topics. Figure 1 illustrates examples of the two types of biases, as well as unbiased stance rationale. Furthermore, our results reveal a significant negative correlation between stance bias and stance detection performance, emphasizing the necessity of alleviating bias to improve the effectiveness of LLMs in stance detection.
Existing research of debiasing in stance detection largely centered on the creation of unbiased training samples and the retraining of stance detection models Kaushal et al. (2021); Yuan et al. (2022b). However, there are two core limitations to the application of these debiasing methods in LLMs. Limitation#1, research Luo et al. (2023) has shown that such retraining processes will undermine the generality of LLMs, potentially leading to catastrophic forgetting; not to mention that there are restrictions with certain closed-source LLMs like GPT-3.5-turbo, which can only be accessed with a restricted inference API, preventing access to internal model parameters. Limitation#2, existing approaches to constructing unbiased training samples typically entail the analysis of prevalent bias patterns, subsequently automating their construction based on these identified patterns, exemplified by substituting ‘Men’ with ‘Women’. However, when dealing with stance detection tasks, our forthcoming analysis illuminates that these samples display varying bias propensities, attributable to divergences in sentiments and stance objectives. Consequently, utilizing conventional methods to create unbiased samples poses a significant challenge.
Therefore, to address the above two limitations, we propose to mitigate biases of LLMs in stance detection with a Counterfactual Augmented Calibration Network, coined as FACTUAL. We establish a trainable calibration network to approximate the inverse projection function of the bias label distribution within LLMs. This calibration network receives samples, along with stance judgments and rationales from LLMs, and generates calibrated stance judgments. We construct counterfactual augmented data against the training data to rectify stance biases. The counterfactual samples are constructed from both causal and non-causal features, which can enhance the calibration network to yield unbiased stances and accomplish out-of-domain generalization. Through counterfactual augmented supervised training, the calibration network can capture biases present in specific samples, thereby performing debiasing. The main contributions of our work are summarized as follows:
1) We are the first to investigate the biases of LLMs on stance detection, categorizing the biases into two main types from the perspective of causality and proposing metrics to quantify these two types of biases.
2) We propose FACTUAL, a novel framework called the Counterfactual Augmented Calibration Network to mitigate biases of LLMs on stance detection.
3) A series of experiments demonstrate that our FACTUAL can effectively reduce the bias of LLMs in stance detection, improving the performance in both in-target and zero-shot stance detection tasks [^3] The code is available at https://github.com/Leon-Francis/FACTUAL.
2 Related Work
Biases in Large Language Models
Some studies Gonçalves and Strubell (2023) have examined the biases existing in Large Language Models (LLMs), these biases mainly include gender and religion Salinas et al. (2023), politics Jenny et al. (2023); He et al. (2023), and spurious correlations Zhou et al. (2023). The associated debiasing efforts are centered around retraining the language model with debiased samples Dong et al. (2023); Limisiewicz et al. (2023). Zheng et al. (2023) found that LLMs are vulnerable to option position changes in MCQs due to their inherent ‘selection bias’. They perform debiasing by approximating the overall bias distribution. while based on our analysis in Section 3, the bias distribution varies significantly across different stance detection samples, so this method is not applicable.
Mitigating Biases in Stance Detection
Currently, studies developed for mitigating biases in stance detection are oriented toward fine-tuned models. Kaushal et al. (2021) analyzed two biases existing in the current datasets: target-independent lexical choices and target-independent sentiment-stance correlations, and built an unbiased dataset. Yuan et al. (2022a) incorporated the stance reasoning process as task knowledge to retrain the model to reduce bias. Yuan et al. (2022b) constructed unbiased samples through counterfactual reasoning and performed adversarial bias learning. These methods involve retraining models and constructing unbiased training samples through special marks, which cannot be directly applied to LLMs.
3 Biases of LLMs in Stance Detection
3.1 Bias Measurement
Stance bias refers to the systematic errors where models tend to choose certain stances due to the influence of specific biases and stereotypes. Inspired by Zheng et al. (2023), the standard deviation of recalls (RStd) on stance labels is an excellent metric for quantitatively measuring systematic errors. The formula is as follows:
$RStd=\sqrt{\frac{1}{K}\sum_{i=1}^{K}\left(\frac{TP_{i}}{P_{i}}-\frac{1}{K}\sum_{j=1}^{K}\frac{TP_{j}}{P_{j}}\right)^{2}}$ (1)
Where $K$ is the number of stance labels, $TP_{i}$ is the number of true positive instances for stance label $i$, and $P_{i}$ is the number of instances of stance label $i$. This measurement resists label imbalance and effectively reflects the model’s bias tendency on samples (Refer to Appendix A for the validation).
3.2 Experimental Result
Through statistical analysis of the results from LLMs, we identified two significant types of bias: Sentiment-stance Spurious Correlations and Target Preference Bias.
3.2.1 Sentiment-Stance Spurious Correlations
Figure 2: The recall score of each stance label on three sentiment subsets, normalizing by subtracting the overall recall scores of the corresponding stance labels across overall dataset, on Sem16, P-Stance, and VAST. POS for positive, NEU for neutral, NEG for negative.
Sentiment can influence stance judgment but is not the primary determinant. Overreliance on sentiment by the model suggests susceptibility to sentiment-stance spurious correlations, leading to biased stance assessments. To investigate stance bias across different sentiments, we first ascertain the sentiment label for each sample. In the Sem16 dataset, each sample has annotated sentiment labels, categorized as positive, neutral, or negative. For the P-Stance and VAST datasets, we utilize GPT-4 to annotate the sentiment labels. To gain a preliminary understanding of sentiment-stance spurious correlations, we first divide the dataset into three subsets based on sentiment categories. For each subset, we calculate the recall score for each stance label and normalize it by subtracting the overall recall score of the corresponding stance labels across the dataset, as shown in Figure 2. Observations indicate that LLMs tend to erroneously predict ‘support’ for positively-sentiment samples and ‘against’ for negatively-sentiment ones, indicating a deviation from expected patterns and highlighting an inherent stance bias. Hence, we identify the Sentiment-stance Spurious Correlations (SSC) as a type of bias in LLMs on stance detection.
We calculate the average of the RStd across all sentiments as our quantification for sentiment-stance spurious correlations:
[…]
where $X_{s}$ represents instances with sentiment label $s$, $|S|$ denotes the number of sentiment labels, which in our experiment, is 3.
We conducted experiments in various settings: Task-Des used task-related descriptions for stance judgment, CoT-Demo used the task description with 4-shot chain-of-thought demonstration, and Debias-Instruct used the task description indicating that sentiment was spurious cues for stance judgment. Refer to Appendix C for the detailed prompts. The results are shown in Table 1. We can observe that in most cases, there is a negative correlation between bias-SSC and stance detection performance. See further analysis in Appendix B. Moreover, prompt engineering methods proved ineffectual in mitigating this inherent bias.
Sem16 P-Stance VAST
[…]
LLaMA-2-70b-chat
Task-Des 17.80 60.08 23.36 79.89 16.87 68.36
CoT-Demo 27.52 58.68 22.81 80.77 22.55 67.08
Debias-Instruct 19.24 63.62 24.86 78.85 19.63 68.68
GPT-3.5-Turbo-0125
Task-Des 27.13 52.82 23.72 81.62 28.70 49.86
CoT-Demo 18.08 67.59 22.75 80.88 16.32 69.90
Debias-Instruct 23.75 51.77 23.48 81.48 30.53 48.68
Table 1: Bias-SSC and macro F1-score of stance detection on the Sem16, P-Stance and VAST dataset. Refer to Appendix D for detailed results on each sentiment.
Figure 3: The recall score of each stance label on several target subsets, normalizing by subtracting the overall recall score of the corresponding stance labels across all targets, on Sem16, P-Stance, and VAST dataset. HC for Hillary Clinton, LA for Legalization of Abortion, AT for Atheism, JB for Joe Biden, BS for Bernie Sanders, DT for Donald Trump, CH for Christian, CL for Election, HP for Humanity Program.
3.2.2 Target Preference Bias
LLMs exhibit bias towards certain individuals or topics. This bias can interfere with their ability to judge stances based on the text, leading to biased stance judgments. We refer to this bias as target preference bias. To preliminarily observe the target preference bias of LLMs, we randomly sampled some targets from different datasets, calculated the recall scores for each stance label on each target subset, and normalized it by subtracting the overall recall score of the corresponding stance labels across all targets, as shown in Figure 3. We observed that, on different targets, LLMs displayed markedly different tendencies in stance selection, which ultimately affected the correctness of stance judgment. Therefore, we identify the Target Preference Bias (TPB) as a type of bias in LLMs on stance detection.
We calculate the average of the RStd of all targets as our quantification for target preference bias:
[…]
where $X_{t}$ represents instances with stance target $t$, $|T|$ denotes the number of targets.
We conduct experiments based on Task-Des, CoT-Demo, and Debias-Instruct which emphasize the need to judge the stance based on the text and not to include the inherent attitude towards the target. Refer to Appendix C for the detailed prompts. The results are shown in Table 2. We can observe that in most cases, there is a negative correlation between bias-TPB and stance detection performance. See further analysis in Appendix B, and prompt engineering fails to effectively mitigate bias-TPB.
Sem16 P-Stance VAST
[…]
LLaMA-2-70b-chat
Task-Des 17.59 60.08 9.09 79.89 7.76 68.36
CoT-Demo 27.56 58.68 11.57 80.77 9.64 67.08
Debias-Instruct 16.37 61.40 8.94 78.70 4.86 69.10
GPT-3.5-Turbo-0125
Task-Des 22.64 52.82 5.43 81.62 28.44 49.86
CoT-Demo 13.47 67.59 6.61 80.88 8.40 69.90
Debias-Instruct 21.87 53.33 5.79 81.59 26.77 51.66
Table 2: Bias-TPB and macro F1-score of stance detection on the Sem16, P-Stance and VAST dataset. Refer to Appendix D for detailed results on each target.
Figure 4: The overall architecture of our proposed FACTUAL. (a) and (b) in the counterfactual data generation represent two ways to generate counterfactual augmentation. $X$ donates the text, $T$ donates the target, $H$ donates the features of the interaction of text and target, and $Y$ donates the stance label. $C$ represents confounding factors, which arise from the two types of biases previously analyzed and may distort the stance prediction.
4 Mitigating Bias with Calibration
Given $\{x_{n},t_{n}\}_{n=1}^{N}$ as the labeled dataset, where $x$ denotes the input text and $t$ denotes the corresponding target, LLMs obtain the stance predictions $y$ through the task instructions $\mathcal{I}$ for stance detection: $P_{obs}(y_{i}|\mathcal{I};x_{i},t_{i})$. Inspired by Zheng et al. (2023), we believe that it can be deconstructed into the unbiased distribution $P_{unbiased}$ of the LLMs performing the stance detection task, and the bias distribution $P_{bias}$ formed by confounding $C_{i}$:
$P_{obs}=P_{unbiased}(y_{i}|x_{i},t_{i})P_{bias}(y_{i}|C_{i})$ (4)
The confounding factor $C_{i}$ arises from the two types of biases analyzed earlier. It is important to note that it may be unaffected, influenced by a single bias, or impacted by both types of biases. Specifically, $P_{obs}$ denotes the stance judgment text, which can be regarded as the outcome of word probability distribution following the argmax operation, derived from LLMs. We aim to estimate the unbiased stance distribution $P_{unbiased}$.
4.1 Calibration Network
By estimating the bias distribution based on the overall distribution of known samples (from the training set), we can obtain unbiased outputs by multiplying the observed distribution of LLMs by the inverse of the approximated bias distribution:
$P_{unbiased}=P_{obs}(y_{i}^{\prime}|\mathcal{I};x_{i},t_{i})\widetilde{P}_{bias}(y_{i}^{\prime\prime}|C_{i})^{-1}$
where $y_{i}^{\prime}$ and $y_{i}^{\prime\prime}$ represent the label distribution output by LLMs and the label distribution affected by bias, $\widetilde{P}_{bias}$ represents the estimate of bias.
However, based on the bias analysis in Section 3, we found that for stance detection, the samples with different sentiments and stance targets have completely different stance bias distributions. Therefore, we propose employing a network to capture the bias distribution specific to each sample, called the calibration network $f_{Cal}$. We use the network $f_{Cal}$ to approximate the inverse projection function of the bias distribution:
$f_{Cal}=P_{bias}(y_{i}^{\prime\prime}|C_{i})^{-1}$ (6)
By inputting the predicted stance distribution $P_{obs}$ from LLMs, an approximating unbiased label can be obtained:
$P_{unbiased}(\hat{y_{i}})=f_{Cal}(P_{obs}(y_{i}^{\prime}|\mathcal{I};x_{i},t_{i}))$ (7)
Specifically, as illustrated in Figure 4, we first use the CoT-Demo instruction (refer to Appendix C for the detail) to obtain the stance judgment and rationale from the LLMs. Then, we input the sample, along with this stance judgment and rationale, into our calibration network (using RoBERTa-base in our setup) to obtain the debiased stance output. We train the calibration network using the cross-entropy loss function with ground truth label:
$\mathcal{L}_{CE}=-\sum_{i=1}^{N}y_{i}\log(f_{Cal}(P_{obs}(y_{i}^{\prime}|\mathcal{I};x_{i},t_{i}))$ (8)
Through supervised training, our calibration network can capture biases present in specific samples, thereby performing debiasing.
4.2 Counterfactual Data Augmentation
One challenge in supervised training is the limited representation of bias within the overall training set, and the learned bias features are difficult to generalize. To fully leverage our analysis of the existing stance biases in LLMs (Section 3), facilitate the calibration network to learn diverse bias patterns, we generate the Counterfactual Augmented Data (CAD) against the training data.
First, we construct counterfactual samples by perturbing non-causal features, approaching from the perspective of bias causation. As shown in counterfactual data generation (a) in Figure 4, to mitigate the impact of sentiment-Stance spurious correlations, we perturb the text $x_{i}$ against the spurious correlation of sentiment. To mitigate the impact of target preference bias, we perturb the target $t_{i}$ against the target preference. Specifically, we construct an instruction that allows the LLMs to rephrase the original sentence and target with different words and sentiments while ensuring that the semantics and the stance remain unchanged. Refer to Figure 10 for the detailed prompts. This obtains the perturbed text $x_{i}^{*}$ and perturbed target $t_{i}^{*}$. Since we only disturbed confounding, the stance label remains unaffected. We construct cross-entropy loss on non-causal counterfactual augmented data as follows:
$\mathcal{L}_{CAD}^{\textit{n-cau}}=-\sum_{i=1}^{N}y_{i}\log(f_{Cal}(P_{obs}(y_{i}^{\prime}|\mathcal{I};x_{i}^{*},t_{i}^{*}))$
Second, we construct counterfactual samples from the perspective of actual stance causality. As shown in counterfactual data generation (b) in Figure 4, we make necessary alterations to text $x_{i}$ to reverse the stance to target $t_{i}$, thereby only perturbing the causal features. Refer to Figure 11 for the detailed prompts. This obtains the perturbed text $\widetilde{x_{i}}$ expressing a reversed stance to target $t_{i}$. We construct cross-entropy loss on causal counterfactual augmented data as follows:
$\mathcal{L}_{CAD}^{\textit{cau}}=\sum_{i=1}^{N}y_{i}\log(f_{Cal}(P_{obs}(y_{i}^{\prime}|\mathcal{I};\widetilde{x_{i}},t_{i})))$
4.3 Training Objective
The final training objective incorporates counterfactual augmented data and performs joint training:
$\mathcal{L}=\mathcal{L}_{CE}+\mathcal{L}_{CAD}^{\textit{n-cau}}+\mathcal{L}_{CAD}^{\textit{cau}}$ (11)
5 Experimental Setup
5.1 Datasets
We conduct experiments of in-target and zero-shot stance detection on three benchmark datasets: SemEval-2016 Task6 (Sem16) Mohammad et al. (2016), P-Stance Li et al. (2021) and Varied Stance Topics (VAST) Allaway and McKeown (2020). The statistic of datasets is shown in Table 3.
Dataset Target Favor Against Neutral
Sem16 HC 163 565 256
FM 268 511 170
LA 167 544 222
A 124 464 145
CC 335 26 203
DT 148 299 260
P-Stance Biden 3217 4079 -
Sanders 3551 2774 -
Trump 3663 4290 -
VAST - 6952 7297 4296
Table 3: Statistics of SemEval-2016 Task6, P-Stance and VAST datasets.
Sem16(%) P-Stance(%)
HC FM LA A CC Avg Biden Sanders Trump Avg
Fine-tuning Based Methods
Roberta 55.97 68.19 67.60 65.40 43.08 58.71 84.29 79.56 82.70 82.18
BERTweet 62.31 64.20 64.14 68.12 41.30 57.99 78.09 81.02 82.48 80.53
KPT 71.30 63.30 63.50 - - - 80.40 77.10 80.20 79.23
KEprompt 77.10 68.30 70.30 - - - 84.40 - 83.20 -
WS-BERT-Dual 75.26 66.02 70.42 71.57 57.31 68.12 83.50 79.00 85.80 82.77
KASD-BERT 77.60 70.38 72.29 72.32 61.47 70.81 85.66 80.39 85.35 83.80
LLaMA-2 Based Methods
LLaMA-2-Task-Des 75.96 66.60 61.68 53.40 73.56 66.24 84.31 77.29 78.08 79.89
LLaMA-2-CoT-Demo 74.84 71.45 62.67 57.58 73.26 67.96 85.03 79.77 77.52 80.77
KASD-LLaMA-2 77.89 67.29 52.00 35.78 47.12 56.02 79.59 71.32 67.89 72.93
LLaMA-2-7b-FT 81.86 71.58 65.56 68.74 75.59 72.67 85.79 81.25 87.47 84.84
[…]
- w/o CAD 78.00 70.82 65.57 71.40 72.24 71.61 85.35 82.00 85.51 84.29
GPT-3.5-Turbo Based Methods
GPT-3.5-Turbo-Task-Des 73.33 66.81 67.22 25.18 72.54 61.02 83.20 80.02 81.66 81.62
GPT-3.5-Turbo-CoT-Demo 81.58 73.42 68.28 64.96 78.35 73.32 83.07 77.98 81.59 80.88
KASD-ChatGPT 80.92 70.37 63.26 61.92 62.72 67.84 84.59 79.96 85.06 83.20
[…]
- w/o CAD 82.38 73.80 63.65 69.21 62.93 70.39 85.40 81.36 85.00 83.92
[…]
5.2 Implementation Details
For GPT-3.5-turbo, we utilize GPT-3.5-turbo-0125. For LLaMA2-70b, we utilize LLaMA2-70b-chat. We set the temperature to 1.0, top p to 1.0, max tokens to 1024, and fixed the decoding seed to ensure the reproducibility of our experiments. For our calibration network, we employ the RoBERTa-base model Liu et al. (2019). For our counterfactual data augmentation, we employ GPT-3.5-turbo-0301 to generate counterfactual samples, guided by the instructions detailed in Appendix C. We use AdamW as an optimizer with a batch size of 32. Learning rate is set to 1e-5 and weight decay is set to 1e-3. All training was conducted on NVIDIA A100 40G GPUs. We report averaged scores of 5 runs to obtain statistically stable results.
5.3 Evaluation Metric
Across three datasets, we used the same evaluation metric established by their proposers, which was also adopted by most of the subsequent baselines. We adopt the macro-average of the F1-score as the evaluation metric. For Sem16 and P-Stance, we report $F1=(F_{favor}+F_{against})/2$. For VAST, we report $F1=(F_{favor}+F_{against}+F_{none})/3$. For in-target stance detection, we select each target to divide training, validation, and test sets. In zero-shot stance detection, for Sem16 and P-Stance, we use the leave-one-target-out evaluation setup. For the VAST dataset, we use their original zero-shot dataset settings. We use standard train/validation/test splits for in-target and zero-shot stance detection across the three datasets.
5.4 Comparison Models
The fine-tuned model baselines include vanilla RoBERTa Liu et al. (2019), domain pre-trained model: BERTweet Nguyen et al. (2020), prompt tuning method KPT Shin et al. (2020), joint contrastive learning framework: JointCL Liang et al. (2022), incorporating ConceptGraph knowledge model: KEprompt Huang et al. (2023), incorporating Wikipedia knowledge model: TarBK-BERT Zhu et al. (2022) and WS-BERT He et al. (2022), incorporating knowledge from LLMs: KASD-BERT Li et al. (2023). For large language models, we compare baselines include Task-Des Zhang et al. (2022), CoT-Demo Zhang et al. (2023b), the self-consistent chain-of-thought: CoT-SC Wang et al. (2023), incorporating Wikipedia knowledge for retrieval-augmented generation: KASD-ChatGPT and KASD-LLaMA-2 Li et al. (2023), fine tuning LLaMA-2-7b using LoRA with training set: LLaMA-2-7b-FT, utilizing collaborative role-infused LLM-based agents: COLA Lan et al. (2023) and utilizing logically consistent chain-of-thought: LC-CoT Zhang et al. (2023a) [^4] Since JointCL, TarBK-BERT, COLA, and LC-CoT were only proposed in zero-shot stance detection scenarios, our comparisons with them are conducted solely within the corresponding experimental settings..
Sem16(%) P-Stance(%) VAST(%)
DT HC FM LA A CC Avg Biden Sanders Trump Avg All
Fine-tuning Based Methods
Roberta 32.12 43.45 40.38 38.79 26.80 18.70 33.37 76.29 72.07 67.56 71.97 73.18
BERTweet 26.88 44.82 21.97 31.91 30.49 12.48 28.09 73.13 68.22 67.66 69.67 71.10
JointCL 50.50 54.80 53.80 49.50 54.50 39.70 50.47 - - - - 72.30
TarBK-BERT 50.80 55.10 53.80 48.70 56.20 39.50 50.68 75.49 70.45 65.80 70.58 73.60
KASD-BERT 54.74 64.78 57.13 51.63 55.97 40.11 54.06 79.04 75.09 70.84 74.99 76.82
LLaMA-2 Based Methods
LLaMA-2-Task-Des 66.03 73.79 71.03 66.00 60.44 61.91 66.53 82.81 78.00 78.87 79.89 68.54
LLaMA-2-CoT-Demo 58.56 72.09 73.83 66.10 57.58 62.47 65.11 83.97 79.26 77.96 80.40 67.28
KASD-LLaMA-2 - 77.70 65.57 57.07 39.55 50.72 - 75.28 74.09 69.27 72.88 43.42
LLaMA-2-7b-FT 63.99 55.49 59.46 33.18 46.37 58.24 52.79 83.93 77.00 74.35 78.43 77.80
[…]
- w/o CAD 61.99 69.22 62.77 60.39 40.83 63.69 59.81 83.09 78.21 76.74 79.35 76.61
GPT-3.5-Turbo Based Methods
GPT-3.5-Turbo-Task-Des 61.72 72.70 71.71 67.89 28.87 59.36 60.38 84.08 80.38 82.38 82.28 50.21
GPT-3.5-Turbo-CoT-Demo 64.16 78.69 73.22 72.84 65.15 75.20 71.54 84.08 80.12 82.24 82.15 70.14
KASD-ChatGPT 64.23 80.32 70.41 62.71 63.95 55.83 66.24 83.60 79.66 84.31 82.52 67.03
COLA 71.20 75.90 69.10 71.00 62.30 64.00 68.92 - - - - 73.40
LC-CoT 71.70 82.90 70.40 63.20 - - - - - - - 72.50
[…]
- w/o CAD 63.28 72.65 60.88 62.07 41.65 67.80 61.39 84.26 77.80 75.26 79.11 77.50
[…]
6 Experimental Results
6.1 In-Target Stance Detection
We perform experiments on Sem16 and P-Stance for in-target stance detection. The results are presented in Table 4. It shows that our FACTUAL outperforms all baselines based on different large language models. We can observe that FACTUAL - w/o CAD, which is without the counterfactual data enhancement, the calibration network trained exclusively on the training set data can still improve stance detection performance. Moreover, the application of counterfactual data enhancement furthers the model’s performance. When compared to the LLaMA-2-7b-FT method, which fine-tunes LLaMA-2-7b, our method attains superior accuracy in stance detection with significantly reduced computation resources.
6.2 Zero-Shot Stance Detection
We conduct experiments on Sem16, P-Stance, and VAST for zero-shot stance detection. The results are shown in Table 5. It shows that our FACTUAL outperforms all baselines including both fine-tuned models and large language models. This indicates that our FACTUAL has strong generalization capabilities and can perform well on unseen targets. We can observe that FACTUAL - w/o CAD exhibits subpar performance. This can be attributed to the constraints posed by the exclusive reliance on fine-tuning within the limited training dataset, making it an uphill task to generalize the model’s debiasing capability. Conversely, employing our counterfactual data enhancement bolsters the out-of-domain generalization prowess of the model considerably, yielding impressive results in zero-shot performance. Compared to the LLaMA-2-7b-FT, which performs poorly on zero-shot tasks, FACTUAL demonstrates strong generalization capabilities.
6.3 Mitigating Biases Effect Analysis
Sem16 P-Stance VAST
[…]
LLaMA-2 Based Methods
Task-Des 17.80 17.59 23.36 9.09 23.87 17.76
CoT-Demo 27.52 27.56 22.81 11.57 22.55 9.64
CoT-SC 33.67 27.18 29.85 6.34 31.98 23.70
LLaMA-2-7b-FT 22.44 25.09 18.14 5.07 22.36 6.84
KASD-LLaMA-2 18.74 10.90 18.74 4.43 20.51 18.00
[…]
- w/o CAD 15.43 12.07 19.15 5.81 21.89 11.42
GPT-3.5-Turbo Based Methods
Task-Des 27.13 22.64 23.72 5.43 28.70 28.44
CoT-Demo 18.08 13.47 22.75 6.61 16.32 18.40
CoT-SC 21.42 16.03 26.30 8.06 22.58 21.63
KASD-ChatGPT 19.49 20.17 20.90 13.94 20.54 15.56
[…]
- w/o CAD 12.92 11.83 18.00 4.38 17.27 12.04
Table 6: Results of Bias-SSC and Bias-TPB in in-target stance detection on Sem16, P-Stance, and zero-shot stance detection on VAST. The best scores are in bold.
We conduct experiments to evaluate the Bias-SSC and Bias-TPB of LLMs and further assess the impact of our bias mitigation efforts. The results are shown in Table 6, which indicate that our FACTUAL can effectively alleviate Bias-SSC and Bias-TPB for both GPT-3.5-turbo and LLaMA2-70b, thus validating its effectiveness in mitigating biases. The inclusion of counterfactual data augmentation can effectively improve its debiasing ability, indicating the importance of our counterfactual data augmentation. Our findings also highlight that the integration of counterfactual data augmentation enhances the debiasing capacity of the model, thereby emphasizing the significance of this augmentation in our methodology.
6.4 Ablation Study
Sem16 VAST
[…]
w/o Calibration 67.96 27.52 27.56 67.28 22.55 9.64
[…]
- w/o non-causal 71.29 13.12 13.25 79.38 24.04 7.12
- w/o causal 69.39 10.66 14.82 77.45 21.69 3.97
[…]
w/o Calibration 73.32 18.08 13.47 70.14 16.32 18.40
[…]
- w/o non-causal 74.08 14.36 11.80 79.38 26.48 9.32
- w/o causal 71.71 10.22 10.21 77.80 22.25 6.81
Table 7: Experimental results of ablation study of in-target stance detection on the Sem16, and zero-shot stance detection on VAST. Results on P-Stance are shown in Table 10. The best scores are in bold.
We conduct ablation studies to examine the impact of different components in our FACTUAL: (1) "w/o Calibration" denotes without the calibration network, letting the LLMs directly output stance labels. (2) "w/o CAD" denotes without the counterfactual augmented data when training the calibration network. (3) "w/o non-causal" denotes without the non-causal counterfactual augmented data when training the calibration network. (4) "w/o causal" denotes without the causal counterfactual augmented data when training the calibration network.
The results are presented in Table 7. Note that despite utilizing the same stance reasoning, a lack of calibration can result in sub-optimal results and notable biases. Thus validating the effectiveness of our gate calibration network. Analysis in Appendix E demonstrate that our method exhibits robustness across different prompt templates.
Additionally, removing non-causal counterfactual data significantly increases bias, highlighting its crucial role in bias mitigation. Conversely, eliminating causal counterfactual data markedly reduces performance, underscoring its substantial impact on the accuracy and generalizability of the calibration network.
7 Conclusion
In this paper, we categorize the biases of LLMs in stance detection into two types from the perspective of causality and propose metrics to quantify these biases. Then, we propose a Counterfactual Augmented Calibration Network, coined as FACTUAL. In which, a trainable calibration network and counterfactual data augmentation are explored to mitigate the biases of LLMs in stance detection. Experimental results on in-target and zero-shot stance detection show that our FACTUAL can effectively reduce the bias of LLMs in stance detection and contribute to improved performance.
Limitations
Our framework involves using GPT-3.5 to generate counterfactual augmented data. As we discussed in Appendix F, these samples may contain errors, but overall are beneficial to the training of our calibration network. The methods of constructing counterfactual augmented data using manual annotation or other methods remain to be explored.
Ethics Statement
The datasets used in this paper are sourced from open-access datasets. The VAST dataset provides complete text data in open access. In compliance with the privacy agreement of Twitter for academic usage, the Sem16 and P-Stance were accessed using the official Twitter API [^5] https://developer.twitter.com/en/docs/twitter-api through the Tweet IDs to fetch complete text data. We removed the information on user privacy from the data. In these datasets, we analyze the biases and stereotypes in stance detection for some sensitive targets (e.g., belief, politics, etc.). We DO NOT critique any biases and stereotypes. We focus on analyzing their impacts on stance detection and mitigating these impacts. We used the counterfactual augmented data obtained from the GPT-3.5-Turbo API service from OpenAI. We followed their term and policies. Some examples in our paper may include a stance or tendency. It should be clarified that they are randomly sampled from the dataset for better studying the dataset and task, and do not represent any personal viewpoints.
Appendix A Bias Measurement Analysis
The standard deviation of recalls (RStd) can largely mitigate the impact of label imbalance in the dataset and effectively assess model bias. This metric is commonly used to evaluate model bias and, as highlighted in relevant research Zheng et al. (2023), "This measurement is intuitive that greater recall imbalance indicates more pronounced selection bias and is not as susceptible to label imbalance as the counting-based measurement."
To illustrate the robustness of this metric against label imbalance, we provide an example from our experiments with the Sem16 dataset. The ground truth data distribution is shown in Table 8. We observed a certain degree of imbalance in label distribution. Based on this, we sampled 200 instances in which ground truth is favor, none, and against, respectively on each sentiment. The sampled instances exhibited no imbalance. We calculated Bias-SSC on these sampled instances, obtaining a result of 27.34, compared to 27.13 on the original dataset. Thus, we believe that the actual distribution of sentiment, target, and stance labels, does not significantly affect our measurement of stance detection bias in LLMs.
(%) Positive Neutral Negative
Favor 9.96 1.99 12.79
None 7.52 2.96 15.32
Against 13.82 1.46 34.19
Table 8: The ground truth data distribution of the Sem16 dataset.
Appendix B Bias Influence
Bias represents a macro-level error pattern. In our experiments, the macroscopic results indicate that larger bias-SSC and bias-TPB tend to lead to poorer stance detection results. To prove this, we conducted a comprehensive analysis involving the calculation of Pearson correlation coefficients between bias-SSC and stance detection F1 scores, as well as bias-TPB and stance detection F1 scores, across 97 groups derived from Tables 1-2671011-15 (with bias-SSC and bias-TPB in Tables 6710 and their corresponding F1 values in Tables 4, 5, and bias-SSC and F1 or bias-TPB and F1 values in Tables 11-15). The results are shown in Table 9, which indicate significant negative correlations for bias-SSC and stance detection F1, and for bias-TPB and stance detection F1.
Bias Type Correlation Coefficient p-value
bias-SSC -0.4047 3.9e-05
bias-TPB -0.5323 2.0e-08
Table 9: Results of Pearson correlation coefficients and p-value between bias-SSC and stance detection F1 scores, as well as bias-TPB and stance detection F1 scores.
Appendix C Prompts Setting
We present the prompt templates used in Section 3.2 and Section 4.2.
Specifically, Figure 5 shows the prompt template we use with GPT-4 to obtain sentiment labels. Figures 67, and 8 show the prompt templates corresponding to the Task-Des, CoT-Demo, and Debias-Instruct prompt settings in Section 3.2.1, respectively. Similarly, Figures 67, and 9 display the prompt templates corresponding to the Task-Des, CoT-Demo, and Debias-Instruct prompt settings in Section 3.2.2, respectively.
Figures 10 and 11 illustrate the prompt templates used to obtain non-causal and causal counterfactual augmented data, respectively, as discussed in Section 4.2. Figure 10 presents the constructed instruction for acquiring non-causal counterfactual augmented data, while Figure 11 shows the instruction for obtaining causal counterfactual augmented data.
Figure 5: Prompt template of sentiment labels annotation by GPT-4. Fill the blue text with the corresponding text from the sample.
Figure 6: Prompt template with Task-Des setting. We first outline the stance detection task, then instruct the LLMs to determine the stance based on the sentence in relation to the target. Fill the blue text with the corresponding text and target from the sample.
Figure 7: Prompt template with CoT-Demo setting. We randomly select 4 samples from the training set, provide the ground truth stance labels, and guide GPT-4 to generate chain-of-thought rationales as examples for this prompt. Fill the green text with constructed examples, and fill the blue text with the corresponding text and target from the sample.
Figure 8: Prompt template with SSC Debias-Instruct setting. We add explicit debiasing instructions following the task description. Fill the blue text with the corresponding text and target from the sample.
Figure 9: Prompt template with TPB Debias-Instruct setting. We add explicit debiasing instructions following the task description. Fill the blue text with the corresponding text and target from the sample.
Figure 10: Prompt template that allows the LLMs to rephrase the original sentence with different words and sentiments and express the target while ensuring that the semantics and the stance of the perturbed sample towards the target remain unchanged. Fill the blue text with the corresponding text, target, and stance label from the sample.
Figure 11: Prompt template that makes necessary modifications to reverse the applicability of the label. Fill the blue text with the corresponding text, target, and stance label from the sample.
Appendix D Experimental Result of LLMs Bias
P-Stance
[…]
w/o Calibration 67.96 27.52 27.56
[…]
- w/o non-causal 71.29 13.12 13.25
- w/o causal 69.39 10.66 14.82
[…]
w/o Calibration 73.32 18.08 13.47
[…]
- w/o non-causal 74.08 14.36 11.80
- w/o causal 71.71 10.22 10.21
Table 10: Experimental results of ablation study of in-target stance detection on the P-Stance dataset. The best scores are in bold.
Table 10 shows the experimental results of the ablation study on the P-Stance dataset, which are consistent with our previous observations. We present the complete experimental results in Section 3. Tables 1112, and 13 show the RStd and the macro F1-Score of samples with different sentiment on the Sem16, P-Stance, and VAST datasets. Tables 14 and 15 present the RStd and the macro F1-Score of samples with different stance target on the Sem16, P-Stance, and VAST datasets. We can observe that in most cases, a larger stance bias leads to poorer stance detection results. In Tables 1112, and 13, samples with positive and negative emotions exhibited larger Rstd, indicating that sentiment influenced the stance judgment of LLMs as a bias pattern. In Tables  14 and 15, on some controversial debate topics such as the "Legalization of Abortion", the "Feminist Movement", and specific individuals like "Donald Trump" and "Hillary Clinton", larger Rstd indicates that LLMs demonstrated a relatively large target preference bias.
Sem16(%)
Positive Neutral Negative
[…]
LLaMA-2-70b-chat
Task-Des 22.20 59.73 23.65 58.11 7.56 65.19
CoT-Demo 26.15 58.24 27.88 55.18 28.54 62.04
Debias-Instruct 26.86 55.04 25.51 58.20 5.34 68.61
GPT-3.5-Turbo-0125
Task-Des 29.37 48.37 25.22 58.78 26.80 55.44
CoT-Demo 16.02 65.77 25.47 61.69 12.74 70.69
Debias-Instruct 30.52 46.09 16.31 60.07 24.43 54.91
Table 11: RStd of sentiment labels and macro F1-score of stance detection on Sem16 dataset. The best scores are in bold.
P-Stance(%)
Positive Neutral Negative
[…]
LLaMA-2-70b-chat
Task-Des 33.32 68.42 19.98 62.68 16.77 72.97
CoT-Demo 32.62 67.17 15.10 65.00 20.70 73.61
Debias-Instruct 34.66 66.49 22.81 59.14 17.10 71.82
GPT-3.5-Turbo-0125
Task-Des 33.79 68.61 23.48 62.32 13.88 75.66
CoT-Demo 32.39 66.99 19.25 65.09 16.60 74.44
Debias-Instruct 33.56 68.00 20.05 63.19 16.83 75.07
Table 12: RStd of sentiment labels and macro F1-score of stance detection on P-Stance dataset. The best scores are in bold.
VAST(%)
Positive Neutral Negative
[…]
LLaMA-2-70b-chat
Task-Des 30.19 61.47 5.79 60.85 14.63 66.97
CoT-Demo 36.20 59.60 18.93 67.87 12.51 63.87
Debias-Instruct 34.60 56.63 10.82 59.49 13.48 67.61
GPT-3.5-Turbo-0125
Task-Des 36.33 44.58 19.75 52.72 30.03 46.74
CoT-Demo 26.01 69.05 8.78 66.73 14.17 67.38
Debias-Instruct 38.40 40.96 23.51 49.01 29.66 46.19
Table 13: RStd of sentiment labels and macro F1-score of stance detection on VAST dataset. The best scores are in bold.
Sem16(%)
HC FM LA A CC
[…]
LLaMA-2-70b-chat
Task-Des 11.59 73.64 24.41 58.33 6.84 59.36 15.12 48.94 29.98 60.15
CoT-Demo 28.36 64.71 35.98 52.96 19.27 57.10 19.91 55.87 34.26 62.78
Debias-Instruct 8.73 76.50 14.75 63.39 9.77 58.64 25.76 39.32 22.83 69.16
GPT-3.5-Turbo-0125
Task-Des 28.68 61.97 24.21 57.47 26.98 56.75 5.81 27.54 27.52 60.36
CoT-Demo 20.99 74.12 15.16 65.61 11.12 64.74 6.75 57.89 13.30 75.57
Debias-Instruct 28.07 63.07 27.81 54.72 24.08 57.11 5.43 29.14 23.94 62.63
Table 14: RStd of targets and macro F1-score of stance detection on Sem16 dataset. The best scores are in bold.
P-Stance(%) VAST(%)
JB BS DT ALL
[…]
LLaMA-2-70b-chat
Task-Des 2.69 84.31 8.79 77.29 15.79 78.08 7.76 68.36
CoT-Demo 7.91 85.03 7.62 79.77 19.19 77.52 9.64 67.08
Debias-Instruct 1.58 82.63 9.29 75.10 15.96 78.36 4.86 69.10
GPT-3.5-Turbo-0125
Task-Des 0.53 83.20 4.82 80.02 10.94 81.66 28.44 49.86
CoT-Demo 3.22 83.07 4.01 77.98 12.59 81.59 8.40 69.90
Debias-Instruct 0.63 82.91 5.36 79.01 11.39 82.85 26.77 51.66
Table 15: RStd of targets and macro F1-score of stance detection on P-Stance and VAST dataset. The best scores are in bold.
Appendix E Prompt Robustness
We conducted experiments, and demonstrating that our method is robust across different prompt templates. The results are shown in Table 16, and the prompts used are listed in Table 17. According to experimental results, the variance in direct stance inference through prompts (w/o Calibration) is 3.5454 for LLaMA-2 and 10.5929 for GPT-3.5-Turbo. In contrast, the variance for stance judgments output by our FACTUAL is 0.3766 for LLaMA-2 and 1.1055 for GPT-3.5-Turbo. We believe this robustness stems from the fact that our calibration network uses LLM-generated rationales to analyze the stance of samples.
Sem16 (LLaMA-2-70b-chat) Sem16 (GPT-3.5-Turbo-0125)
Avg Bias-SSC Bias-TPB Avg Bias-SSC Bias-TPB
FACTUAL-Prompt-1 73.82 9.61 5.52 76.16 11.31 7.74
w/o Calibrition 67.96 27.52 27.56 73.32 18.08 13.47
w/o CAD 71.61 15.43 12.07 70.39 12.92 11.83
[…]
- w/o CAD-causal 69.39 10.66 14.82 71.71 10.22 10.21
FACTUAL-Prompt-2 72.48 9.99 7.69 74.73 10.86 8.77
w/o Calibrition 66.97 26.92 23.75 69.50 22.22 15.88
w/o CAD 70.51 19.26 13.58 70.49 19.89 12.40
[…]
- w/o CAD-causal 69.89 10.03 11.57 70.81 11.58 10.56
FACTUAL-Prompt-3 73.74 9.19 5.60 73.59 10.11 7.77
w/o Calibrition 67.37 19.45 18.14 65.35 26.35 25.88
w/o CAD 69.99 12.83 13.15 67.86 16.84 18.24
[…]
- w/o CAD-causal 69.90 8.11 10.27 69.60 9.53 13.16
Table 16: Experimental results of three different prompt templates on in-target stance detection on the Sem16 dataset. The best scores are in bold.
Task Description Stance detection is to determine the attitude or tendency towards a certain target through a given sentence, including favor, against, and neutral. **Please read the following examples carefully and use them as references to judge the attitude of the sentence towards the target.**
[…]
Table 17: Three different prompts used in the experiment in Table 16.
Appendix F Human Evaluation of Counterfactual Augmented Data
CAD-non-causal CAD-causal
Quality Achievement Quality Achievement
Human Eval 96.67% 91.33% 95.33% 88.83%
Table 18: Human evaluation on Sem16, P-Stance, and VAST datasets. Quality stands for qualitative assessment of the generated samples and Achievement stands for the achievement of generating objectives.
We randomly select 500 samples and use human evaluation (with three experienced researchers who are not involved in this work and have worked on natural language processing for over 3 years) to measure the counterfactual data generated by GPT-3.5-turbo. The primary consideration focuses on the qualitative assessment of the generated samples, necessitating evaluators to confirm the accuracy of both the grammar and the affirmed stance. The secondary consideration pertains to achieving generating objectives, necessitating evaluators to confirm if the samples were generated as guideline instructions. Evaluators respond to these considerations with a binary "yes" or "no". Subsequently, we calculate the average ratio of affirmative responses from three evaluators for each query. The results in Table 18 show that the generated samples are of high quality, contributing substantially to our calibration network training.
Appendix G Case Study
We conduct a case study on Sem16, P-Stance and VAST datasets, to analyze the biases of LLMs in the stance detection task and the practical effectiveness of our calibration network. The results are show in Table 19, 20 and 21. The correct analysis patterns of LLMs are marked in blue, while biased analysis patterns are marked in red. We can observe that for some samples with strong sentiment expressions, such as the examples in Table 20, LLMs are influenced by sentiment Spurious cues and result in biased stance judgments. For some controversial debate topics, such as the examples in Table 19, LLMs generate hallucinations due to their preferences, leading to biased stance judgments.
Text (Target: Legalization of Abortion) Label LLMs Pred LLMs Stance Rationale Calibration
A prochoice advocate but circumcise ur baby? Fucking hypocrite! #circumcision #humanrights Against Favor The sentence criticizes the hypocrisy of being a prochoice advocate but supporting circumcision, indicating a negative attitude towards circumcising babies. Since the author is prochoice, it can be inferred that they are in favor of the legalization of abortion. Against
Text (Target: Feminist Movement) Label LLMs Pred LLMs Stance Rationale Calibration
Equality is the police burying a domestic violence accusation against a female sports star, too #wedidit #usa Against Favor The sentence mentions burying a domestic violence accusation against a female sports star, which implies a negative view towards this unequal treatment. Therefore, the attitude of the sentence towards the Feminist Movement is likely favorable. Against
Table 19: Case study on the Sem16 dataset.
Text (Target: Donald Trump) Label LLMs Pred LLMs Stance Rationale Calibration
You might as well put a sock in it. Voters see these coordinated stupid attacks are like your #Kavanaugh symphony. #Dems can’t win with hollow attempts to bash #Trump. Favor Against The sentence uses derogatory terms like ’stupid attacks’ and ’hollow attempts to bash #Trump,’ indicating a critical perspective. Therefore, the attitude of the sentence toward ’Donald Trump’ is against. Favor
Text (Target: Joe Biden) Label LLMs Pred LLMs Stance Rationale Calibration
Saying that is as absurd as saying breast cancer screening increases breast cancer. Is there any Republican who isn’t as dumb as dirt? Favor Against The sentence implies negative assumptions about the intelligence of Republicans. The mention of Joe Biden in the context of criticizing Republicans suggests a critical or skeptical attitude towards Joe Biden. Favor
Table 20: Case study on the P-Stance dataset.
Text (Target: Homosexual) Label LLMs Pred LLMs Stance Rationale Calibration
Yet that same Bible gives us Christians a particular anthropology, and reveals to us the meaning of sex and marriage within the cosmic order. And herein lies the problem. If you look to a "holy" book written long, long ago and betraying its human origins on every page for guidance on issues of cosmology, biology, sexuality and other human and natural sciences, you will inevitably be in constant conflict with social progress and, indeed, reality. There is no secular opposition to homosexuality." Favor Against The sentence criticizing the use of a ’holy’ book for guidance on issues related to sexuality and natural sciences. It suggests that relying on such sources will lead to conflict with social progress and reality. Additionally, the statement asserts that there is no secular opposition to homosexuality, implying a negative view towards those who oppose it based on religious beliefs. Overall, the tone of the sentence is against homosexuality, indicating a critical stance. Favor
Text (Target: Gun Control) Label LLMs Pred LLMs Stance Rationale Calibration
Military and police, yes. They’re sworn and trained to uphold the country and the laws of the country. The rest is an ego trip. Favor Against The sentence expresses a negative attitude towards gun control in the military and police, stating that they are sworn and trained to uphold the country and its laws. However, it is critical of the use of guns by others, referring to it as an ego trip. Therefore, the attitude of the sentence towards gun control is against Favor
Table 21: Case study on the VAST dataset.